# Decentralized Reliability Estimation for Mixnets

### v1

@@ 3.1 Link performance estimation
Routing of packets. Gateways send packets to the first-layer mix node specified as first hop of in each packet’s route. Mix nodes execute the packet processing algorithm and further relay packets to their next destination. As part of replay protection in mixnets, mix nodes must record a tag per processed packet (denoted by $t$, cf. Sect. 3.2). A newly received packet is then checked against the list of tags already seen: if the packet is fresh, it is processed and its tag is added to the list; and otherwise the packet is dropped. Each node $j$ uses a Bloom filter [2] to store in a space-efficient manner the $\sum_{i}{s}_{(i,j)}$ tags received from its predecessors $i\in P(j)$ during an epoch. When adding a fresh packet tag to the Bloom filter nodes first append a binary flag, indicating whether the packet integrity checks where correct, allowing the packet to be further relayed; or incorrect, forcing node $j$ to drop the packet. Mix node decryption keys are updated per epoch to provide forward security in addition to ensuring that packets cannot be replayed in a later epoch.
Measurement packets have as destination a randomly selected gateway with the probability weights being part of the mixnet’s routing policy. When gateways receive a measurement packet, they store its tag in a locally maintained Bloom filter and proceed to discard the packet. Note that gateways do not need to store in the Bloom filter the tags of non-measurement packets, meaning that their Bloom filters can be smaller than those of mix nodes.
Post-epoch stage: revelation of measurements. Once an epoch has ended, the protocol proceeds in three steps:
@@
    Each gateway $g$ broadcasts the number of packets $N_{g,c}$ sent per credential $c$.
@@
    All nodes and gateways broadcast their Bloom filters.
@@
    Gateways reveal the openings of the ${s}^{*}_{(c,g)}(\approx p_{\mathsf{lot}}\cdot{s}_{(c,g)})$ measurement packets sent from each credential $c$ they have serviced in the epoch.
@@
    Any entity reading broadcast outputs can verify the validity of the gateway openings and use them to recreate the $\sum_{g}\sum_{c}{s}^{*}_{(c,g)}$ measurement packets generated in the epoch. As result, the path and recipient of each measurement packet, as well as the tags seen at each step of the route, become available. All entities can then check which measurement packet tags were included in the Bloom filters broadcast by mix nodes and gateways in step (1).
Given a link $(i,j)$ in the mixnet, we denote by ${s}_{(i,j)}^{*}$ the amount of measurement packets successfully transmitted in the link, and compute it by counting the measurement packets whose (per-hop) tags appear in the Bloom filters of both $i$ and $j$. On the other hand, the quantity ${d}_{(i,j)}^{*}$ denotes the number of measurement packets dropped in the link, computed by counting the measurement packets whose tags appear in the Bloom filter of $i$ but not in the one of $j$. For every intermediary node $j$ it holds that packets received from all predecessors $i$ are either successfully transmitted to their next destination $k$, or dropped before reaching $k$:
$\sum_{i}{s}^{*}_{(i,j)}=\sum_{k}{s}^{*}_{(j,k)}+{d}^{*}_{(j,k)}$ (3)
In the case of gateways, measurement packets revealed by $g$ that do not appear in the Bloom filter of the corresponding first layer node $i$ are considered as part of ${d}^{*}_{(g,i)}$. Similarly, measurement packets addressed to $g$ from last layer node $k$ that are reported by $k$ but do not appear in $g$’s Bloom filter, are considered part of ${d}^{*}_{(k,g)}$.
Packets reported as dropped due to failed integrity checks (as indicated by the flag stored with the tag) are entirely removed from the measurement sample set. Similarly, packets with ‘holes’ in their reported paths, i.e., included in node $j$’s Bloom filter but missing from its predecessor, are also entirely discarded as measurements.
Taking into account that mixnet intermediaries cannot distinguish measurement from regular packets while routing, measurement packets are transmitted and dropped by nodes at the same rate as any other packet. We use this to produce an estimated performance $\hat{\rho}_{e}$ for link $e$, given by the transmitted and dropped samples ${s}_{e}^{*}$ and ${d}_{e}^{*}$:
@@
To bound the estimation error $\epsilon_{e}$ we use methods that estimate coin bias from a sequence of observed coin flips, considering that the coin flip can result in transmission (heads) with probability ${\rho}_{e}$; or in a drop (tails) with probability $1-{\rho}_{e}$. Given an edge with ${s}_{e}^{*}+{d}_{e}^{*}$ observed coin flips and confidence level $Z$, the maximum estimation error $\epsilon_{e}$ is given by:
$\epsilon_{e}=Z\cdot\sqrt{\frac{{\rho}_{e}\cdot(1-{\rho}_{e})}{{s}_{e}^{*}+{d}_{e}^{*}}}\approx Z\cdot\sqrt{\frac{\hat{\rho}_{e}\cdot(1-\hat{\rho}_{e})}{{s}_{e}^{*}+{d}_{e}^{*}}}$ (6)
The maximum estimation error occurs when ${\rho}_{e}=0.5$, and thus $\epsilon_{e}\leq\frac{Z}{2\sqrt{{s}_{e}^{*}+{d}_{e}^{*}}}$ for any value of ${\rho}_{e}$, with the error diminishing significantly when ${\rho}_{e}$ approaches zero or one. Since ${\rho}_{e}$ is not available, in practice we use $\hat{\rho}_{e}$ when computing $\epsilon_{e}$ as best available approximation of ${\rho}_{e}$. In cases where $\hat{\rho}_{e}=1$ or $\hat{\rho}_{e}=0$, we use Laplace’s rule of succession to estimate $\epsilon_{e}\approx\frac{1}{{s}_{e}^{*}+{d}_{e}^{*}+2}$.
Note that the estimation accuracy for edge $e$ is dependent on the number of measurement samples ${s}_{e}^{*}+{d}_{e}^{*}$ but not on the total number of packets ${s}_{e}+{d}_{e}$. The method thus requires an amount of measurement samples per link to have a certain accuracy, which stays constant regardless of the amount of total traffic. Thus, as more traffic is routed in the network, the percentage $p_{\mathsf{lot}}$ of measurement packets can be decreased while maintaining accuracy. The output of this protocol is a value $\hat{\rho}_{e}\approx{\rho}_{e}$ per edge that estimates the link performance, defined as the fraction of successfully transmitted packets, together with a corresponding sampling error $\epsilon_{e}$ that qualifies the accuracy of the estimation for a given confidence level.
@@
$k$: security parameter; length of the hash function output. $p_{\mathsf{lot}}$: probbability of a measurement packet. $ctr$: packet counter $r_{\mathsf{type}}$: randomness for measurement lottery. $r_{\mathsf{pkt}}$: randomness for packet encoding. $y_{n}$: public-key of node $n$ $T$: target that determines measurement lottery outcome $s_{i}$: Diffie Helllman key corresponding to packet hop $i$ $b_{i}$: blinding value corresponding to packet hop $i$ $\tilde{r}_{i}$: encoding randomness for packet hop $i$ $n_{i}$: node corresponding to packet hop $i$ $\mathsf{n}$: unpredictable random nonce drawn from public source $\alpha_{i}$: base element for packet hop $i$ $\psi_{i}$: ciphertext for packet hop $i$ $x_{i}$: secret-key of processing node in packet hop $i$ $t_{i}$: pseudorandom tag calculated in packet hop $i$
@@ 3.2 VRF-Based Routing
In this section we show how to transition our protocol to the full threat model that considers gateways to be utility maximizers while clients may be malicious. Specifically, we consider gateway utility to be a monotonically increasing function [^4] Nym’s reward function [10] can easily be configured for gateways to increase monotonically with all these variables. of (i) link performance, as measured by the protocol itself, for all the mixnet links that gateways are a counterparty to, (ii) client satisfaction, a ratio between the packets that an honest client successfully sends and are delivered divided by the total number of packets the client requests to be sent, (iii) credential consumption, the total number of announced legitimate packets sent, and (iv) not getting caught cheating in any of the requested public operations. The gateway being a utility maximizer means that it in all cases will avoid strategies that reduce its utility but may choose any strategy that maximizes its utility.
Viewed in this setting, the scheme of the previous section can be easily seen to be ineffective. For instance, given that measurement packets are not subject to integrity checks, it is possible for a gateway to follow the protocol as described, but then open as measurement packets some other packets that only partially match the sent measurements. In that case, the performance in the links where the gateway is a counterparty will be correct but performance estimations for other links will be lower, given that incorrect tags are revealed for some hops and thus it seems that the measurements got dropped further in their path. At the same time, since clients can be malicious we cannot rely on them entirely for generating the measurement traffic. To mitigate these issues, we introduce VRF-based routing, where the measurement packets are provably stochastically generated across an epoch’s packet transmissions, emulating independent coin flips.
@@
Second, the processing of a packet $(\alpha_{0},\psi_{0})$ corresponding to VRF values $r_{\mathsf{pkt}},r_{\mathsf{type}}$ proceeds as follows. The gateway calculates $s_{0}=\alpha_{0}^{x_{0}},\tilde{r}_{0}=H(\mathsf{rnd},s_{0})$, where $x_{0}$ is its secret-key and it verifies that: (I) the mixnet encoding was correctly applied to produce $\psi_{0}$ given randomness $\tilde{r}_{0}$ and the public key $y_{0}$; and (II) the per packet randomization key is correctly computed: in the case of a non measurement packet it is $\alpha_{0}=\alpha^{r_{\mathsf{pkt}}}$, while it is $\alpha_{0}=g^{r_{\mathsf{pkt}}}$ in the case of a measurement packet. If these checks pass, it calculates the next hop as $n_{1}=\mathsf{Routing}(n_{0},r_{0})$, where $r_{0}=H(\mathsf{next},s_{0})$ and then it forwards $(\alpha_{1},\psi_{1})$ to node $n_{1}$, where $\alpha_{1}=\alpha_{0}^{b_{0}}$ with $b_{0}=H(\mathsf{bli},s_{0})$ and $\psi_{1}$ is the payload of packet $\psi_{0}$ after processing it in the way the mixnet decoding requires.
@@
Next we illustrate how processing works for the remaining nodes in the path. Node $n_{i}$, given $(\alpha_{i},\psi_{i})$ for $i=1,\ldots,\nu$ operates as follows. First, it processes the packet to reveal payload $\psi_{i+1}$ and if $i<\nu$, the next hop $n_{i+1}$. Then it calculates $s_{i}=\alpha_{i}^{x_{i}}$, $\tilde{r}_{i}=H(\mathsf{rnd},s_{i})$, and if $i<\nu$, $\alpha_{i+1}=\alpha_{i}^{b_{i}}$ where $b_{i}=H(\mathsf{bli},s_{i})$. The node then verifies that (I) the mixnet encoding was correctly applied to produce $\psi_{i}$ given randomness $\rho_{i}$, (II) in case $i<\nu-1$ $n_{i+1}=\mathsf{Routing}(n_{i},r_{i})$ where $r_{i}=H(\mathsf{next},s_{i})$, (the randomized routing check). It also calculates the tag $t_{i}=H(\mathsf{tag},s_{i})$, and stores $(b_{i},\mathrm{flag}_{i})$ in a database of processed tags where $\mathrm{flag}_{i}$ is $1$ if and only if the integrity check of $\psi_{i}$ is valid. Finally, if $i<\nu$, it forwards $(\alpha_{i+1},\psi_{i+1})$ to node $n_{i+1}$. Finally, in case $i=\nu$, the payload $\psi_{\nu+1}$ is parsed as the plaintext and is processed by node $n_{\nu}$.
@@
The key point of the above construction is that the randomness used for encoding is pseudorandomly determined by the key $\alpha$ and the VRF function. Specifically, for each $ctr$, there is a per packet randomization key $\alpha_{0}=g^{x\cdot x_{0}\cdot r_{\mathsf{pkt}}}$ that is used to seed the mixnet encoding. For measurement packets, observe that the gateway can completely validate their computation due to the fact that $r_{\mathsf{pkt}}$ is known and the secret randomization key $x$ of the client is not used in the packet calculations. The same holds true for any party that obtains the value $r_{\mathsf{pkt}}$. Furthermore, any packet can be demonstrated to be a measurement or not, by checking the inequality $r_{\mathsf{type}}<T$ which is true iff the $ctr$ value corresponds to a measurement packet.
@@
The following theorem is a straightforward observation regarding the construction. The first and third statement follow easily from the Diffie Hellman assumption, while the second takes advantage of the random oracle and the fact that routing choices depend on the unpredictable value $\mathsf{nonce}$. Due to lack of space a proof is sketched in the appendix.
@@
Assuming the pseudorandomness of the VRF, the Decisional Diffie Hellman assumption and that $H(\cdot)$ is modeled as a random oracle, (I) the VRF-based encoding presented above does not interfere with the security or the privacy of the underlying mixnet encoding scheme. (II) the packet routes generated by the above process are sampled according to $\mathsf{Routing}()$ pseudorandomly, even in the setting where $vk,\alpha$ are adversarially chosen. (III) assuming client and gateway are honest, the measurement packets are computationally indistinguishable to non-measurement packets.
Deploying VRF-based routing. The concept of VRF-based routing is deployed utilizing (I) a threshold compact e-cash scheme [31], (II) a VRF function based on Dodis Yampolskiy’s PRF [14]; specifically, $\mathsf{K}$ produces a secret key $sk_{\mathsf{vrf}}$ and public-key $vk_{\mathsf{vrf}}=g^{sk_{\mathsf{vrf}}}$, while the VRF evaluation is set to $\mathsf{E}(\mathsf{n}||ctr||\mathrm{label})=(r_{\mathsf{label}},\pi)$ where $r_{\mathsf{label}}=H(vk_{\mathsf{vrf}},g,u_{\mathrm{label}})$, $u_{\mathrm{label}}=g^{1/(sk_{\mathsf{vrf}}+2^{\ell}\mathsf{n}+2\cdot ctr+\mathrm{label})}$, and $\pi_{\mathrm{label}}=(u_{\mathrm{label}},\pi^{\prime}_{\mathrm{label}})$ where and $\pi^{\prime}_{\mathrm{label}}=\mathrm{NIZK}(g,vk_{\mathsf{vrf}},u_{\mathrm{label}},ctr,\ell,\mathsf{nonce},\allowbreak\mathrm{label}:\exists sk_{\mathsf{vrf}}(u_{\mathrm{label}}^{sk_{\mathsf{vrf}}+2^{\ell}\mathsf{nonce}+4ctr+\mathrm{label}+1}=g\land vk_{\mathsf{vrf}}=g^{sk_{\mathsf{vrf}}}))$. Note in the above $\ell,\mathrm{label}\in\mathbb{N},$ are public values, $ctr<2^{\ell-2}$ always where $ctr$ is the packet counter, $\mathsf{nonce}$ is a public random value that is produced after all other public values are committed. Recall that we use the numerical labels $\mathrm{pkt}=0,\mathrm{type}=1,\mathrm{exit}=2$.
@@
In order to engage in VRF routing, the client acquires a long term key $vk_{\mathsf{long}},sk_{\mathsf{long}}$ associated with a compact e-cash wallet that can accommodate $N$ unlinkable payments. We refer to this action as acquiring a “subscription” with $N$ credentials. Subsequently the user registers with a gateway of her choice spending a credential and shares a key $\alpha$ that will be bound to the specific credential. The gateway, in turn, receives an allowance of packets to enter the mixnet equal to the sum of all credentials it receives in each period. To facilitate packet transmission, it registers a VRF key $vk_{\mathsf{vrf}}$ associated with each client’s credential it serves. Importantly note that, before registering each credential, the client is free to switch gateways without being tracked (due to the unlikability of the compact e-cash scheme). At the same time, if the client is found to double spend a credential, this is traced back to her subscription $vk_{\mathsf{long}}$ that can be revoked, hence penalizing the misbehaving client.
@@
Opening a measurement packet. A gateway and/or client can open a measurement packet corresponding to a credential containing $\alpha$ by revealing the values $(ctr,r_{\mathsf{pkt}},\pi_{\mathsf{pkt}},r_{\mathsf{type}},\pi_{\mathsf{exit}})$. In such case anyone can verify that this is a measurement packet corresponding to a counter $ctr$ by checking that $r_{\mathsf{type}}<T$ and the proofs. Moreover, the whole computation of such packet is possible to be verified as it can be derived from $r_{\mathsf{pkt}}$. Observe that the packet is deterministically generated based on the values $ctr,\mathsf{n},r_{\mathsf{pkt}}$. Regarding the size of the opening, observe we need one index, one group element, one hash and two exponents, for each of the three VRF openings. Using Ed25519, this amounts to 388 bytes per measurement packet opening.

### v2

@@ 3.1 Link reliability estimation
Routing of packets. Gateways send packets to the first-layer mix node specified as first hop in each packet’s route. Mix nodes execute the packet processing algorithm and further relay packets to their next destination. As part of replay protection in mixnets, mix nodes must record a tag per processed packet (denoted by $t$, cf. Sect. 3.2). A newly received packet is then checked against the list of tags already seen: if the packet is fresh, it is processed and its tag is added to the list; and otherwise the packet is dropped. Each node $j$ stores in a space-efficient manner the $\sum_{i}{s}_{(i,j)}$ tags received from its predecessors $i\in P(j)$ during an epoch (this can be done by a Bloom filter [2] or a Merkle tree). We call this the “tag commitment” — note that it only needs to be binding and space efficient. When adding a fresh packet tag, the node appends also a binary flag that indicates whether the packet integrity checks were correct. If correct, the packet is allowed to be further relayed. If incorrect, the node drops the packet. Mix node decryption keys may be updated per epoch to provide forward security in addition to ensuring that packets cannot be replayed in a later epoch. Measurement packets have as destination a randomly selected gateway with the probability weights being part of the mixnet’s routing policy. When gateways receive a measurement packet, they store its tag in their tag-commitment and proceed to discard the packet.
Post-epoch stage: revelation of measurements. Once an epoch has ended, the protocol proceeds in four steps:
@@
    Each gateway $g$ broadcasts the number of packets ${s}_{(c,g)}\leq S_{c}$ sent per credential $c$ used in the epoch.
@@
    All mix nodes and gateways broadcast their space efficient tag-commitments.
@@
    Each gateway $g$ broadcasts the openings of the ${s}^{*}_{(c,g)}(\approx p_{\mathsf{lot}}\cdot{s}_{(c,g)})$ measurement packets generated in the epoch, which is followed by broadcasting the openings of the corresponding tag-commitments by all nodes (in case Bloom filters are used for tag-commitments this last step is not necessary).
@@
    Any entity reading broadcast outputs can verify the validity of the gateway openings and use them to recreate the $\sum_{g}\sum_{c}{s}^{*}_{(c,g)}$ measurement packets generated in the epoch. As result, the path and recipient of each measurement packet, as well as the tags seen at each step of the route, become available. Any entity can then check which measurement packet tags were included in the Bloom filters broadcast by nodes in step (2).
Given a link $(i,j)$ in the mixnet, we denote by ${s}_{(i,j)}^{*}$ the amount of measurement packets successfully transmitted in the link, and compute it by counting the measurement packets whose (per-hop) tags appear in the Bloom filters of both $i$ and $j$. On the other hand, the quantity ${d}_{(i,j)}^{*}$ denotes the number of measurement packets dropped in the link, computed by counting the measurement packets whose tags appear in the tag-commitment of $i$ but not in the one of $j$. For every intermediary node $j$ it holds that packets received from all predecessors $i$ are either successfully transmitted to their next destination $k$, or dropped before reaching $k$:
$\sum_{i}{s}^{*}_{(i,j)}=\sum_{k}({s}^{*}_{(j,k)}+{d}^{*}_{(j,k)})$ (3)
In the case of gateways, measurement packets revealed by $g$ that do not appear in the tag-commitment of the corresponding first layer node $i$ are considered as part of ${d}^{*}_{(g,i)}$. Similarly, measurement packets addressed to $g$ from last layer node $k$ that are reported by $k$ but do not appear in $g$’s tag-commitment, are considered part of ${d}^{*}_{(k,g)}$.
Packets reported as dropped due to failed integrity checks (as indicated by the flag stored with the tag) are entirely removed from the measurement sample set. Similarly, packets with ‘holes’ in their reported paths, i.e., included in node $j$’s tag-commitment but missing from its predecessor, are also entirely discarded as measurements.
Taking into account that mixnet intermediaries cannot distinguish measurement from regular packets while routing, measurement packets are transmitted and dropped by nodes at the same rate as any other packet. We use this to produce an estimated reliability $\hat{\rho}_{e}$ for link $e$, given by the transmitted and dropped samples ${s}_{e}^{*}$ and ${d}_{e}^{*}$:
@@
To bound the estimation error $\epsilon_{e}$ we use methods that estimate coin bias from a sequence of observed coin flips, considering that the coin flip can result in transmission (heads) with probability ${\rho}_{e}$; or in a drop (tails) with probability $1-{\rho}_{e}$. Without having any prior for the distribution ${\rho}_{e}$, i.e., assuming all values are equally likely, ${\rho}_{e}$ follows a Beta distribution with density function ${\rho}_{e}^{s^{*}_{e}}(1-{\rho}_{e})^{d^{*}_{e}}/B(s^{*}_{e}+1,d^{*}_{e}+1)$, mean $\hat{\rho}_{e}$ and variance $\hat{\rho}_{e}\cdot(1-\hat{\rho}_{e})/(s^{*}_{e}+d^{*}_{e}+1)$.
In case we have a large number of measurements ${s}_{e}^{*}+{d}_{e}^{*}$ and the coin bias is not very skewed, one can use a normal approximation that yields the Wald interval for the maximum estimation error $\epsilon_{e}$ at confidence level $Z$:
$\epsilon_{e}=Z\cdot\sqrt{\frac{\hat{\rho}_{e}\cdot(1-\hat{\rho}_{e})}{{s}_{e}^{*}+{d}_{e}^{*}}}$ (6)
Note that the above error estimation is inaccurate as $\rho_{e}$ approaches zero or one; an accurate confidence interval can be derived following Clopper Pearson [6] (but it has no nice closed formulation, as the above).
Note that the estimation accuracy for edge $e$ is dependent on the number of measurement samples ${s}_{e}^{*}+{d}_{e}^{*}$ but not on the total number of packets ${s}_{e}+{d}_{e}$. The method thus requires an amount of measurement samples per link to have a certain accuracy, which stays constant regardless of the amount of total traffic. Thus, as more traffic is routed in the network, the percentage $p_{\mathsf{lot}}$ of measurement packets can be decreased while maintaining accuracy. The output of this protocol is a value $\hat{\rho}_{e}\approx{\rho}_{e}$ per edge that estimates the link reliability, defined as the fraction of successfully transmitted packets, together with a corresponding sampling error $\epsilon_{e}$ that qualifies the accuracy of the estimation for a given confidence level.
Communication complexity. For $t$ measurement packets, and assuming the number of layers of the mix-net $L$ is constant (in practice $L=3$), the scheme requires $O(t\log t)$ broadcasts when using Merkle trees for tag-commitments.
@@
$k$: security parameter; length of the hash function output $p_{\mathsf{lot}}$: probability of a measurement packet $ctr$: packet counter $r_{\mathsf{type}}$: randomness for measurement lottery $r_{\mathsf{pkt}}$: randomness for packet encoding $y_{n}$: public-key of node $n$ $T$: target that determines measurement lottery outcome $s_{i}$: Diffie Hellman key corresponding to packet hop $i$ $b_{i}$: blinding value corresponding to packet hop $i$ $\tilde{r}_{i}$: encoding randomness for packet hop $i$ $n_{i}$: node corresponding to packet hop $i$ $\mathsf{n}$: unpredictable random nonce drawn from public source $\alpha_{i}$: base element for packet hop $i$ $\psi_{i}$: ciphertext for packet hop $i$ $x_{i}$: secret-key of processing node in packet hop $i$ $t_{i}$: pseudorandom tag calculated in packet hop $i$
@@ 3.2 VRF-Based Routing
In this section we show how to transition our protocol to the full threat model where clients can be malicious and gateways are utility maximizers. Specifically, we consider gateway utility to be a monotonically increasing function [^2] Nym’s reward function [12] can easily be configured for gateways to increase monotonically with all these variables. of (i) link reliability, as measured by the protocol itself, for all the mixnet links that gateways are a counterparty to, (ii) client satisfaction, a ratio between the packets that an honest client successfully sends and are delivered divided by the total number of packets the client requests to be sent, (iii) credential consumption, the total number of announced legitimate packets sent, and (iv) not getting caught cheating in any of the requested public operations. The gateway being a utility maximizer means that it will avoid strategies that reduce its utility but otherwise engage in any strategy.
Viewed in this setting, the scheme of the previous section can be easily seen to be ineffective. For instance, given that measurement packets are not subject to integrity checks, it is possible for a gateway to follow the protocol as described, but then open as measurement packets some crafted packets that partially match the sent measurements. In that case, the reliability in the links where the gateway is a counterparty will be correct but reliability estimations for other links will be lower, given that incorrect tags are revealed for some hops and thus it seems that the measurements got dropped further in their path. At the same time, since clients can be malicious we cannot rely on them entirely for generating the measurement traffic. To mitigate these issues, we introduce VRF-based routing, where the measurement packets are provably stochastically generated across an epoch’s packet transmissions, emulating independent coin flips.
@@
Second, the processing of a packet $(\alpha_{0},\psi_{0})$ corresponding to VRF values $r_{\mathsf{pkt}},r_{\mathsf{type}},r_{\mathsf{exit}}$ proceeds as follows. The gateway calculates $s_{0}=\alpha_{0}^{x_{0}},\tilde{r}_{0}=H(\mathsf{rnd},s_{0})$, where $x_{0}$ is its secret-key and it verifies that: (I) the mixnet encoding was correctly applied to produce $\psi_{0}$ given randomness $\tilde{r}_{0}$ and the public key $y_{0}$; and (II) the per packet randomization key is correctly computed: in the case of a non measurement packet it is $\alpha_{0}=\alpha^{r_{\mathsf{pkt}}}$, while it is $\alpha_{0}=g^{r_{\mathsf{pkt}}}$ in the case of a measurement packet. If these checks pass, it calculates the next hop as $n_{1}=\mathsf{Routing}(n_{0},r_{0})$, where $r_{0}=H(\mathsf{next},s_{0})$ and then it forwards $(\alpha_{1},\psi_{1})$ to node $n_{1}$, where $\alpha_{1}=\alpha_{0}^{b_{0}}$ with $b_{0}=H(\mathsf{bli},s_{0})$ and $\psi_{1}$ is the payload of packet $\psi_{0}$ after processing it in the way the mixnet decoding requires.
@@
Next we illustrate how processing works for the remaining nodes in the path. Node $n_{i}$, given $(\alpha_{i},\psi_{i})$ for $i=1,\ldots,\nu$, operates as follows. First, it processes the packet to reveal payload $\psi_{i+1}$ and if $i<\nu$, the next hop $n_{i+1}$. Then it calculates $s_{i}=\alpha_{i}^{x_{i}}$, $\tilde{r}_{i}=H(\mathsf{rnd},s_{i})$, and if $i<\nu$, $\alpha_{i+1}=\alpha_{i}^{b_{i}}$ where $b_{i}=H(\mathsf{bli},s_{i})$. The node then verifies that (I) the mixnet encoding was correctly applied to produce $\psi_{i}$ given randomness $\rho_{i}$, (II) in case $i<\nu-1$ $n_{i+1}=\mathsf{Routing}(n_{i},r_{i})$ where $r_{i}=H(\mathsf{next},s_{i})$, (the randomized routing check). It also calculates the tag $t_{i}=H(\mathsf{tag},s_{i})$, and stores $(b_{i},\mathrm{flag}_{i})$ in a database of processed tags where $\mathrm{flag}_{i}$ is $1$ if and only if the integrity check of $\psi_{i}$ is valid. Finally, if $i<\nu$, it forwards $(\alpha_{i+1},\psi_{i+1})$ to node $n_{i+1}$. Finally, in case $i=\nu$, the payload $\psi_{\nu+1}$ is parsed as the plaintext and is processed by node $n_{\nu}$.
@@
The key point of the above construction is that the randomness used for encoding is pseudorandomly determined by the key $\alpha$ and the VRF function. Specifically, for each $ctr$, there is a per packet randomization key $\alpha_{0}=g^{x\cdot r_{\mathsf{pkt}}}$ that is used to seed the mixnet encoding. For measurement packets, observe that the gateway can completely validate their computation due to the fact that $r_{\mathsf{pkt}}$ is known and the secret randomization key $x$ of the client is not used in the packet calculations. The same holds true for any party that obtains the value $r_{\mathsf{pkt}}$. Furthermore, any packet can be demonstrated to be a measurement or not, by checking the inequality $r_{\mathsf{type}}<T$ which is true iff the $ctr$ value corresponds to a measurement packet.
@@
The following theorem is a straightforward observation regarding the construction. The first and third statement follow easily from the Diffie Hellman assumption and the security of the VRF, while the second takes advantage of the random oracle and the fact that routing choices depend on the unpredictable value $\mathsf{nonce}$. A proof is sketched in Appendix 1 and the nonce may be generated from public randomness such as a blockchain [22].
@@
Assuming the pseudorandomness of the VRF, the Decisional Diffie Hellman assumption and that $H(\cdot)$ is modeled as a random oracle. (I) any security or privacy attack against the VRF-based encoding presented above can be transformed to an attack against the underlying mixnet encoding scheme with negligible loss in success probability, (II) the packet routes generated by the above process are sampled according to $\mathsf{Routing}()$ pseudorandomly, even in the setting where $vk,\alpha$ are adversarially chosen. (III) the measurement packets are computationally indistinguishable from non-measurement packets in the view of any relay.
Deploying VRF-based routing. The concept of VRF-based routing is deployed utilizing (I) a threshold compact e-cash scheme [33], (II) a VRF function based on Dodis Yampolskiy’s PRF [18]; specifically, $\mathsf{K}$ produces a secret key $sk_{\mathsf{vrf}}$ and public-key $vk_{\mathsf{vrf}}=g^{sk_{\mathsf{vrf}}}$, while the VRF evaluation is set to $\mathsf{E}(\mathsf{n}||ctr||\mathrm{label})=(r_{\mathsf{label}},\pi)$ where $r_{\mathsf{label}}=H(vk_{\mathsf{vrf}},g,u_{\mathrm{label}})$, $u_{\mathrm{label}}=g^{1/(sk_{\mathsf{vrf}}+2^{\ell}\mathsf{n}+2\cdot ctr+\mathrm{label})}$, and $\pi_{\mathrm{label}}=(u_{\mathrm{label}},\pi^{\prime}_{\mathrm{label}})$ where and $\pi^{\prime}_{\mathrm{label}}=\mathrm{NIZK}(g,vk_{\mathsf{vrf}},u_{\mathrm{label}},ctr,\ell,\mathsf{nonce},\allowbreak\mathrm{label}:\exists sk_{\mathsf{vrf}}(u_{\mathrm{label}}^{sk_{\mathsf{vrf}}+2^{\ell}\mathsf{nonce}+4ctr+\mathrm{label}+1}=g\land vk_{\mathsf{vrf}}=g^{sk_{\mathsf{vrf}}}))$. Note in the above $\ell,\mathrm{label}\in\mathbb{N},$ are public values, $ctr<2^{\ell-2}$ always where $ctr$ is the packet counter, $\mathsf{nonce}$ is a public random value that is generated by the randomness beacon after all other public values are committed on the blockchain. Recall that we use the numerical labels $\mathrm{pkt}=0,\mathrm{type}=1,\mathrm{exit}=2$.
@@
In order to engage in VRF routing, the client acquires a long term key $vk_{\mathsf{long}},sk_{\mathsf{long}}$ associated with a wallet that can accommodate $N$ unlinkable anonymous credentials. We refer to this action as acquiring a “subscription” with $N$ credentials. Subsequently the user registers with a gateway of her choice spending a credential and shares a key $\alpha$ that will be bound to the specific credential. The gateway, in turn, receives an allowance of packets to enter the mixnet equal to the sum of all credentials it receives in each period. To facilitate packet transmission, it registers a VRF key $vk_{\mathsf{vrf}}$ associated with each client’s credential it serves. Importantly note that, before registering each credential, the client is free to switch gateways without being tracked (due to the unlikability of anonymous credentials). At the same time, if the client is found to double spend a credential, this is traced back to her subscription $vk_{\mathsf{long}}$ that can be revoked, hence penalizing the misbehaving client.
@@
Opening a measurement packet. A gateway and/or client can open a measurement packet corresponding to a credential containing $\alpha$ by revealing the values $(ctr,r_{\mathsf{pkt}},\pi_{\mathsf{pkt}},r_{\mathsf{type}},\pi_{\mathsf{type}},r_{\mathsf{exit}},\pi_{\mathsf{exit}})$. In such case anyone can verify that this is a measurement packet corresponding to a counter $ctr$ by checking that $r_{\mathsf{type}}<T$ and the proofs. Moreover, the whole computation of such packet is possible to be verified as it can be derived from $r_{\mathsf{pkt}}$ and $r_{\mathsf{exit}}$. Observe that the packet is deterministically generated based on the values $ctr,\mathsf{n},r_{\mathsf{pkt}},r_{\mathsf{exit}}$. Regarding the size of the opening, observe we need one index, one group element, one hash and two exponents, for each of the three VRF openings. Using Ed25519, this amounts to 388 bytes per measurement packet opening.